# MCGrad: Multicalibration at Web Scale

### v2

@@ Abstract.
We propose MCGrad, a novel and scalable multicalibration algorithm. Multicalibration - calibration in sub-groups of the data - is an important property for the performance of machine learning-based systems. Existing multicalibration methods have thus far received limited traction in industry. We argue that this is because existing methods (1) require such subgroups to be manually specified, which ML practitioners often struggle with, (2) are not scalable, or (3) may harm other notions of model performance such as log loss and Area Under the Precision-Recall Curve (PRAUC). MCGrad does not require explicit specification of protected groups, is scalable, and often improves other ML evaluation metrics instead of harming them. MCGrad has been in production at Meta, and is now part of hundreds of production models. We present results from these deployments as well as results on public datasets.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
A machine learning model is said to be calibrated when its predictions match the true outcome frequencies (Silva Filho et al., 2023; Kull et al., 2017; Bella et al., 2010). The importance of calibration to ML-based systems has been widely recognized (Kendall and Gal, 2017; Guo et al., 2017; Kull et al., 2019), including in web applications such as content ranking (Kweon et al., 2022; Penha and Hauff, 2021), recommender systems (Gulsoy et al., 2025; da Silva and Jannach, 2025; Steck, 2018), digital advertising (Chen and Qi, 2022; Chaudhuri et al., 2017; Fan et al., 2023) and content moderation (Liu et al., 2025; Vidgen et al., 2020; Kivlichan et al., 2021). Calibration is necessary for such systems to make optimal decisions without under- or overestimating risk or opportunity.
@@
Despite its potential, multicalibration has received limited traction in industry. We argue that this is due to three main reasons. First, existing multicalibration methods require protected groups to be manually defined (Jin et al., 2025): not only do users need to specify the covariates that require protection (e.g., user age and user country), but also they must define concrete protected groups through binary membership indicators, such as “is an adult in the US?”. This poses a significant challenge to real-world application of multicalibration because practitioners (a) may not have established the precise set of protected groups, (b) may be subject changes to the definition of protected groups over time thus requiring ongoing effort and change management (e.g., due to changes related to legal context, policy frameworks, ethical considerations, or otherwise), or (c) may only seek an overall performance improvement without prioritizing specific groups for protection. The difficulties defining protected groups become even more pronounced when the set of features that we want to protect is large. Second, existing methods lack the ability to scale to large datasets and/or many protected groups, which makes them hard to deploy in production. For instance, Hebert-Johnson et al. (2018)’s algorithm scales at least linearly in time and memory with the number of groups, being potentially inefficient when a large set of groups is specified. To be deployed at web scale, multicalibration methods must be highly optimized so as to not introduce significant computational requirements at training or inference time. Third, existing multicalibration methods lack guardrails to be safely deployed, and often introduce a risk that multicalibrating a predictor ends up harming model performance may prevent deployment. This may be caused by overfitting.
@@
In this paper, we introduce MCGrad (MultiCalibration Gradient Boosting) a novel multicalibration algorithm deployed in production that is lightweight and safe to use. MCGrad requires only the specification of a set of protected features, rather than predefined protected groups. It then identifies miscalibrated regions within this feature space, and calibrates a base classifier’s predictions over all groups that can be defined based on these features. On a high level, MCGrad uses a Gradient Boosted Decision Tree (GBDT) algorithm recursively, such that the result converges to a multicalibrated predictor (Section 3.1). By using a highly optimized GBDT implementation, like LightGBM (Ke et al., 2017), MCGrad inherits scalability and regularization (Section 3.2). Finally, it employs an early stopping procedure that guarantees, with high probability, that it improves the base predictor’s likelihood and does not overfit, ensuring that model performance is not harmed (Section 3.3).
@@
Empirically, on benchmark datasets, we show that MCGrad improves the base predictor’s outputs in terms of multicalibration, calibration, and even predictive performance (Section 4). We complement this analysis by describing the impact of MCGrad on real-world data: MCGrad has been used in production at Meta on hundreds of machine learning models, and in total generates over a million multicalibrated real-time predictions per second (Section 5). To the best of our knowledge, this is the large-scale adoption of multicalibration in production. With these real-world results, we add to the growing body of evidence that shows that multicalibration can significantly contribute to ML model performance. We also provide valuable practical learnings from applying multicalibration in industry (Section 6), and link multicalibration to related areas (Section 7).
@@ 2. Background on (Multi)Calibration
This gives rise to two ways of interpreting the metric. The absolute scale $\text{ECCE}(f)$ is designed to measure the magnitude of miscalibration. When the data size increases, $n\rightarrow\infty$, $\text{ECCE}(f)$ converges to a true magnitude of miscalibration (0 if $f$ is calibrated and a positive value otherwise). The sigma scale $\nicefrac{{\text{ECCE}(f)}}{{\sigma(f)}}$ however converges to a universal distribution when $f$ is calibrated, and diverges to infinity when it is miscalibrated. The two scales are used to answer different questions. ECCE is used to answer the question ”is the miscalibration large?”. $\nicefrac{{\text{ECCE}(f)}}{{\sigma(f)}}$ is used to answer the question ”is there statistical evidence of miscalibration?”. For example, we might have that $\text{ECCE}(f_{1})=0.05=20\sigma(f_{1})$ and $\text{ECCE}(f_{2})=0.05=2\sigma(f_{2})$. This would tell us that while the magnitude of miscalibration on the data appears to be similar for $f_{1},f_{2}$, the result is statistically significant for $f_{1}$ but not for $f_{2}$ could simply be due to random sampling. See (Arrieta-Ibarra et al., 2022; Tygert, 2023) for further discussion.
@@
Multicalibration is defined in various ways across the literature (Hebert-Johnson et al., 2018; Shabat et al., 2020; Błasiok et al., 2024; Gopalan et al., 2022b; Deng et al., 2023; Haghtalab et al., 2023a). Here, we employ an operational definition of multicalibration:
@@
A common choice is to use a uniform bound on the MC-deviation, which is accommodated by choosing the scale parameter $\tau_{h}(f)$ to be a constant equal to $1$. In Appendix A.1 we show that Definition 2.1, with $\tau_{h}(f)^{2}=\mathbb{E}\left[h(X)f(X)(1-f(X))\right]$, naturally matches with Multicalibration Error (MCE) by Guy et al. (2025), which quantifies multicalibration error by computing the maximum ECCE over the protected groups with a suitable normalization:
@@
Using a multicalibration metric that is equivalent to the definition of $\alpha$-MC can help ensure that empirical work is aligned with theory. In Section 4 we use MCE to measure multicalibration, thereby directly estimating the minimal $\alpha$ for which a model is $\alpha$-MC. To our knowledge, this is the first empirical multicalibration study that achieves this. See Appendix A.1 for further discussion.
@@ 3. MCGrad: A Practical Algorithm for Multicalibration
Our proposed algorithm MCGrad, presented in Algorithm 1, is a lightweight algorithm that multicalibrates a base model without requiring the specification of the groups on which the model has to be calibrated. MCGrad relies on a simple key observation: by including the base model’s predictions as a feature, the loss function of the multicalibration algorithm captures the feature values that correspond to miscalibrated initial predictions. In addition, decreasing the loss function implies correcting the base model’s predictions for some regions of the feature space. Because corrections in some regions may negatively impact others, MCGrad runs multiple rounds, using the previous round’s processed predictions as input. Since GBDTs are regularized using shrinkage (or a step size <1<1), a simple rescaling step reduces the number of trees and rounds required, while having a negligible impact on overfitting. Finally, MCGrad employs early stopping to avoid overfitting.
@@ 3.1. Achieving Multicalibration with $T$ Rounds
Our key intuition builds on the following two key insights. First, GBDT returns a solution $f$ that approximately sets the gradient of the loss function $\mathcal{L}$ to zero:
@@
Intuitively, if for some probabilistic predictor $f^{*}$, $f_{T}\rightarrow f^{*}$ as $T\rightarrow\infty$, then for all $h\in\mathcal{H}^{*}$, $\mathbb{E}_{D}\left[h(X,f_{T-1}(X))(Y-f_{T}(X))\right]\rightarrow\mathbb{E}_{D}\left[h(X,f^{*}(X))(Y-f^{*}(X))\right]=0$ as $T\rightarrow\infty$, which is our desired condition. This suggests that, for sufficiently large $T$, the probabilistic predictor $f_{T}$ should be approximately multicalibrated. See Section A.2 for further intuitions and theoretical results.
@@ 3.2. Fast Training and Prediction
To be deployed at web scale there are strict requirements for an algorithm to be efficient both at training and inference. Several design choices ensure that MCGrad meets these requirements.
@@
First, the algorithm has been designed to rely on a relatively small number of calls to a GBDT, delegating the most compute intensive steps to one of many highly optimized GBDT implementations (e.g.  (Ke et al., 2017; Chen and Guestrin, 2016)). This differentiates it from existing implementations such as (Hebert-Johnson et al., 2018). In our implentation, we use LightGBM.
@@
GBDTs are regularized in multiple ways, including using a step size smaller than $1$ for scaling each additional tree. While this is helps to avoid overfitting, it results in a predictor which can be strictly improved by rescaling it by a factor slightly greater than $1$. As a result, the next round will attempt to apply this rescaling. Since linear rescaling is not easily expressed by decision trees, this may require many trees, ending up requiring unnecessarily many trees to achieve multicalibration.
@@ 3.3. Preventing Overfitting
While the multiple rounds are required for convergence, they also increase the capacity of the model. With $T$ rounds and $M$ trees in each round, the MCGrad model is at least as expressive as a tree ensemble with $T\cdot M$ trees. Since the model capacity of tree ensembles is $\Omega(\text{number of trees})$ (see (Shalev-Shwartz and Ben-David, 2013)), overfitting is more likely for a large number of rounds.
We solve this problem with a standard early stopping procedure on the number of rounds $T$. Specifically, we split i.i.d. the dataset $D$ into training ($D_{train}$) and validation ($D_{valid}$) sets, and determine the optimal number of rounds $T$ by taking the last round before the expected loss over the validation set increases. That is,
@@
where the models $f_{t}$ are obtained using $D_{train}$ instead of $D$. Using early stopping in a real-world deployed system has a relevant consequence: with high probability, MCGrad does not harm the base model’s prediction, and, instead, would select $T=0$ as optimal number of rounds if the first step decreased the initial performance:
@@ 4. Benchmark Experimental Analysis
We evaluate MCGrad on both public and production data. In this section, we present experiments on benchmark datasets and compare MCGrad against state-of-the-art baselines. In Section 5 we present results of MCGrad in production. Here, we focus on four research questions:
@@ 4.1. Experimental Setup
We use a total of $11$ tabular benchmark datasets. Five of the datasets are used in  Hansen et al. (2024)’s evaluation of multicalibration methods. We add datasets by using alternative definitions of the ground truth, using other available columns in the American Community Survey (ACS) data. We describe the datasets in more detail in Appendix B. The datasets range from sample size 11k Medical Expenditure Panel Survey (MEPS) to 3.2M (ACSEmploy), and from 10 features (ACSIncome) to 139 (MEPS).
@@
1) Predictive performance. We quantify using Area Under the Precision-Recall curve (PRAUC) (Davis and Goadrich, 2006) and log loss.
@@
3) Multicalibration. Błasiok et al. (2024) proposed the maximum group-wise smECE metric, which was used for empirical evaluations in Hansen et al. (2024). Guy et al. (2025) propose an extension of ECCE to quantify multicalibration, called Multicalibration Error (MCE), which we defined in Eq. (2). Maximum group-wise smECE and MCE have in common that they both are defined as a maximum of some quantity over groups, but differ in the calibration quantity calculated in each group, where they respectively use smECE and ECCE. In this paper, we use MCE to quantify multicalibration. In Appendix B we additionally report group-wise smECE, thereby providing results that are in-line with (Hansen et al., 2024). We can measure any multicalibration metric with respect to either of two sets of groups: Prespecified Groups: refers to a small (less than $15$ across all datasets) set of protected groups that are formed using various features of the datasets. For this, we use the groups specifications that were used in  Hansen et al. (2024). Unspecified Groups: covers the case where the user does not specify which groups to protect against. In that case, the scores are expected to be multicalibrated with respect to all possible groups, and to calculate this, an extensive set of combinatorially generated groups are used.
@@
MCGrad uses the same hyperparameters as in production deployment (see Appendix C). For HKRR we follow Hansen et al. (2024) and pick the best hyperparameter from a set of four specifications using a held-out validation set. For DFMC we use the default hyperparameters of LightGBM since the paper does not suggest any specific hyperparameters.

### v3

@@ Abstract.
We propose MCGrad, a novel and scalable multicalibration algorithm. Multicalibration - calibration in subgroups of the data - is an important property for the performance of machine learning-based systems. Existing multicalibration methods have thus far received limited traction in industry. We argue that this is because existing methods (1) require such subgroups to be manually specified, which ML practitioners often struggle with, (2) are not scalable, or (3) may harm other notions of model performance such as log loss and Area Under the Precision-Recall Curve (PRAUC). MCGrad does not require explicit specification of protected groups, is scalable, and often improves other ML evaluation metrics instead of harming them. MCGrad has been in production at Meta, and is now part of hundreds of production models. We present results from these deployments as well as results on public datasets. We provide an open source implementation of MCGrad at https://github.com/facebookincubator/MCGrad.
@@ 1. Introduction
A machine learning model is said to be calibrated when its predictions match the true outcome frequencies (Silva Filho et al., 2023; Kull et al., 2017; Bella et al., 2010). The importance of calibration to ML-based systems has been widely recognized (Kendall and Gal, 2017; Guo et al., 2017; Kull et al., 2019), including in web applications such as content ranking (Kweon et al., 2022; Penha and Hauff, 2021), recommender systems (Gulsoy et al., 2025; da Silva and Jannach, 2025; Steck, 2018), digital advertising (Chen and Qi, 2022; Chaudhuri et al., 2017; Fan et al., 2023) and content moderation (Liu et al., 2024; Vidgen et al., 2020; Kivlichan et al., 2021). Calibration is necessary for such systems to make optimal decisions without under- or overestimating risk or opportunity.
@@
Despite its potential, multicalibration has received limited traction in industry. We argue that this is due to three main reasons. First, existing multicalibration methods require protected groups to be manually defined (Jin et al., 2025): not only do users need to specify the covariates that require protection (e.g., user age and user country), but also they must define concrete protected groups through binary membership indicators, such as “is an adult in the US?”. This poses a significant challenge to real-world application of multicalibration because practitioners (a) may not have established the precise set of protected groups, (b) may be subject to changes to the definition of protected groups over time thus requiring ongoing effort and change management (e.g., due to changes related to legal context, policy frameworks, ethical considerations, or otherwise), or (c) may only seek an overall performance improvement without prioritizing specific groups for protection. The difficulties defining protected groups become even more pronounced when the set of features that we want to protect is large. Second, existing methods lack the ability to scale to large datasets and/or many protected groups, which makes them hard to deploy in production. For instance, Hebert-Johnson et al. (2018)’s algorithm scales at least linearly in time and memory with the number of groups, being potentially inefficient when a large set of groups is specified. To be deployed at web scale, multicalibration methods must be highly optimized so as to not introduce significant computational requirements at training or inference time. Third, existing multicalibration methods lack guardrails to be safely deployed, and risk harming model performance (e.g., due to overfitting), which may prevent practitioners from deploying them altogether.
@@
In this paper, we introduce MCGrad (MultiCalibration Gradient Boosting) a novel multicalibration algorithm deployed in production that is lightweight and safe to use. MCGrad requires only the specification of a set of protected features, rather than predefined protected groups. It then identifies miscalibrated regions within this feature space, and calibrates a base classifier’s predictions over all groups that can be defined based on these features. On a high level, MCGrad uses a Gradient Boosted Decision Tree (GBDT) algorithm recursively, such that the result converges to a multicalibrated predictor (Section 3.1). By using a highly optimized GBDT implementation, like LightGBM (Ke et al., 2017), MCGrad inherits scalability and regularization (Section 3.2). Finally, it employs an early stopping procedure to avoid overfitting, ensuring that model performance is not harmed (Section 3.3).
@@
Empirically, on benchmark datasets, we show that MCGrad improves the base predictor’s outputs in terms of multicalibration, calibration, and even predictive performance (Section 4). We complement this analysis by describing the impact of MCGrad on real-world data: MCGrad has been used in production at Meta on hundreds of machine learning models, and in total, generates over a million multicalibrated real-time predictions per second (Section 5). MCGrad successfully addresses all three previous challenges and, to the best of our knowledge, this is the largest-scale adoption of multicalibration in production. Our deployment results show significant impact: on $24$ of $27$ models tested via A/B testing on our Looper platform (Markov et al., 2022), MCGrad significantly outperformed Platt scaling, leading to promotion of the MCGrad-calibrated variants to production. Additionally, across 120120+ models on our internal ML platform, MCGrad improved log loss for $88.7\%$ of models, PRAUC for $76.7\%$, and Expected Calibration Error for $86.0\%$. With these real-world results, we add to the growing body of evidence that shows that multicalibration can significantly contribute to model performance. We also provide valuable practical learnings from applying multicalibration in industry (Section 6), and link multicalibration to related areas (Section 7).
@@ 2. Background on (Multi)Calibration
This gives rise to two ways of interpreting the metric. The absolute scale $\text{ECCE}(f)$ is designed to measure the magnitude of miscalibration. When the data size increases, $n\rightarrow\infty$, $\text{ECCE}(f)$ converges to a true magnitude of miscalibration (0 if $f$ is calibrated and a positive value otherwise). The sigma scale $\nicefrac{{\text{ECCE}(f)}}{{\sigma(f)}}$ however converges to a distribution when $f$ is calibrated, and diverges to infinity when it is miscalibrated. The two scales are used to answer different questions. ECCE is used to answer the question ”is the miscalibration large?”. $\nicefrac{{\text{ECCE}(f)}}{{\sigma(f)}}$ is used to answer the question ”is there statistical evidence of miscalibration?”. For example, we might have that $\text{ECCE}(f_{1})=0.05=20\sigma(f_{1})$ and $\text{ECCE}(f_{2})=0.05=2\sigma(f_{2})$. This would tell us that while the magnitude of miscalibration on the data appears to be similar for $f_{1},f_{2}$, the result is statistically significant for $f_{1}$ but not for $f_{2}$, which could simply be due to random sampling. See (Arrieta-Ibarra et al., 2022; Tygert, 2023) for further discussion.
@@
Multicalibration is defined in various ways across the literature (Hebert-Johnson et al., 2018; Shabat et al., 2020; Błasiok et al., 2024; Gopalan et al., 2022b; Deng et al., 2023; Haghtalab et al., 2023). Here, we employ an operational definition of multicalibration:
@@
A common choice is to use a uniform bound on the MC-deviation, which is accommodated by choosing the scale parameter $\tau_{h}(f)$ to be a constant equal to $1$. In Appendix A, we show that Definition 2.1, with $\tau_{h}(f)^{2}=\mathbb{E}\left[h(X)f(X)(1-f(X))\right]$, naturally matches with Multicalibration Error (MCE) by Guy et al. (2025), which quantifies multicalibration error by computing the maximum ECCE over the protected groups with a suitable normalization:
@@
Using a multicalibration metric that is equivalent to the definition of $\alpha$-MC can help ensure that empirical work is aligned with theory. In Section 4 we use MCE to measure multicalibration, thereby directly estimating the minimal $\alpha$ for which a model is $\alpha$-MC. To our knowledge, this is the first empirical multicalibration study that achieves this. See Appendix A for further discussion.
@@ 3. MCGrad: A Practical Algorithm for Multicalibration
Our proposed algorithm MCGrad, presented in Algorithm 1, is a lightweight algorithm that multicalibrates a base model without requiring the specification of the groups on which the model has to be calibrated. MCGrad relies on a key observation: by including the base model’s predictions as a feature, the loss function of the multicalibration algorithm captures the feature values that correspond to miscalibrated initial predictions. In addition, decreasing the loss function implies correcting the base model’s predictions for some regions of the feature space. Because corrections in some regions may negatively impact others, MCGrad runs multiple rounds, using the previous round’s processed predictions as input. Since GBDTs are regularized using shrinkage (or a step size <1<1), a simple rescaling step reduces the number of trees and rounds required, while having a negligible impact on overfitting. Finally, MCGrad employs early stopping to avoid overfitting.
@@ 3.1. Achieving Multicalibration with $T$ Rounds
Our intuition builds on the following two key insights. First, GBDT returns a solution $f$ that approximately sets the gradient of the loss function $\mathcal{L}$ to zero:
@@
Intuitively, if for some probabilistic predictor $f^{*}$, $f_{T}\rightarrow f^{*}$ as $T\rightarrow\infty$, then for all $h\in\mathcal{H}^{*}$, $\mathbb{E}_{D}\left[h(X,f_{T-1}(X))(Y-f_{T}(X))\right]\rightarrow\mathbb{E}_{D}\left[h(X,f^{*}(X))(Y-f^{*}(X))\right]=0$ as $T\rightarrow\infty$, which is our desired condition. This suggests that, for sufficiently large $T$, the probabilistic predictor $f_{T}$ should be approximately multicalibrated. See Appendix B for further intuitions and theoretical results.
@@ 3.2. Fast Training and Prediction
To be deployed at web scale, there are strict requirements for an algorithm to be efficient both at training and inference time. Several design choices ensure that MCGrad meets these requirements.
@@
First, the algorithm has been designed to rely on a relatively small number of calls to a GBDT, delegating the most compute intensive steps to one of many highly optimized GBDT implementations (e.g.  (Ke et al., 2017; Chen and Guestrin, 2016)). This differentiates it from existing implementations such as (Hebert-Johnson et al., 2018). In our implementation, we use LightGBM.
@@
GBDTs are regularized in multiple ways, including using a step size smaller than $1$ for scaling each additional tree. While this helps to avoid overfitting, it results in a predictor which can be improved by rescaling it by a factor slightly greater than $1$. As a result, the next round will attempt to apply this rescaling. Since linear rescaling is not easily expressed by decision trees, this may require many trees, ending up requiring unnecessarily many trees to achieve multicalibration.
@@ 3.3. Preventing Overfitting
While multiple rounds are required for convergence, they also increase the capacity of the model. With $T$ rounds and $M$ trees in each round, the MCGrad model is at least as expressive as a tree ensemble with $T\cdot M$ trees. Since the model capacity of tree ensembles is $\Omega(\text{number of trees})$ (see (Shalev-Shwartz and Ben-David, 2014)), overfitting is more likely for a large number of rounds.
We solve this problem with a standard early stopping procedure on the number of rounds $T$. Specifically, we split i.i.d. the dataset $D$ into training ($D_{train}$) and validation ($D_{valid}$) sets, and determine the number of rounds $T$ by taking the last round before the expected loss over the validation set increases. That is,
@@
where the models $f_{t}$ are obtained using $D_{train}$ instead of $D$. Using early stopping in a real-world deployed system has a relevant consequence: MCGrad does not harm the base model’s prediction, and, instead, would select $T=0$ as optimal number of rounds if the first step decreased the initial performance:
@@ 4. Benchmark Experimental Analysis
We evaluate MCGrad on both public and production data. In this section, we present experiments on benchmark datasets and compare MCGrad against state-of-the-art baselines. In Section 5 we present results of MCGrad in production. Here, we focus on five research questions: [^2] Code: https://github.com/facebookresearch/mcgrad_multicalibration_at_web_scale.
@@
  • Q5.
    How does MCGrad’s computational time compare to existing multicalibration methods?
Dataset # Samples $n$ # Features $d$ Class Distr. Protected Attributes
MEPS (Sharma et al., 2021) 11079 139 0.169 Race, Gender, Age, Income, Insurance Status
Credit (Yeh, 2009) 30000 118 0.221 Gender, Age, Education, Marital Status
Marketing (Moro et al., 2014) 45211 41 0.116 Age, Marital Status, Education, Occupation
HMDA (Cooper et al., 2023) 114185 89 0.752 Race, Gender, Ethnicity, Age
ACSIncomeCA (Ding et al., 2021; Flood et al., 2021) 195665 10 0.410 Race, Gender, Age, Education, Income, Employment
ACSMobility (Ding et al., 2021; Flood et al., 2021) 616207 23 0.735 Race, Gender, Age, Education, Income, Employment
ACSPublic (Ding et al., 2021; Flood et al., 2021) 1123374 24 0.293 Race, Gender, Age, Education, Income, Employment
ACSTravel (Ding et al., 2021; Flood et al., 2021) 1458542 19 0.437 Race, Gender, Age, Education, Income, Employment
ACSIncome (Ding et al., 2021; Flood et al., 2021) 1655429 10 0.370 Race, Gender, Age, Education, Income, Employment
ACSEmploy (Ding et al., 2021; Flood et al., 2021) 3207990 20 0.456 Race, Gender, Age, Education, Income, Employment
ACSHealth (Ding et al., 2021; Flood et al., 2021) 3207990 31 0.150 Race, Gender, Age, Education, Income, Employment
Table 1. Number of samples, features, class distribution ($p(Y=1)$), and protected attributes for each benchmark dataset.
@@ 4.1. Experimental Setup
Our experimental analysis uses 11 datasets (see Table 1). Six binarized prediction tasks are derived from the American Community Survey (ACS) (Ding et al., 2021; Flood et al., 2021) using the folktable package, plus a California-specific income task (ACSIncomeCA). The remaining five datasets are: UCI Bank Marketing (Moro et al., 2014) (term deposit prediction), UCI Default of Credit Card Clients (Yeh, 2009) (debt default), Home Mortgage Disclosure Act (HMDA) (Cooper et al., 2023) (mortgage acceptance), and Medical Expenditure Panel Survey (MEPS) (Sharma et al., 2021) (medical visits). For all datasets we use the same protected groups as in (Hansen et al., 2024). Protected attributes are selected following standard fairness benchmarks.
@@
1) Predictive performance. We employ the Area Under Precision-Recall Curve (PRAUC) (Davis and Goadrich, 2006) and the log loss.
@@
3) Multicalibration. Błasiok et al. (2024) proposed the maximum group-wise smECE metric, which was used for empirical evaluations in Hansen et al. (2024). Guy et al. (2025) proposed an extension of ECCE to quantify multicalibration, called Multicalibration Error (MCE), which we defined in Eq. (2). Maximum group-wise smECE and MCE have in common that they both are defined as a maximum of some quantity over groups, but differ in the calibration quantity calculated in each group, where they respectively use smECE and ECCE. In this paper, we use MCE to quantify multicalibration. In the online supplement [^4] https://github.com/facebookresearch/mcgrad_multicalibration_at_web_scale, we additionally report group-wise smECE, thereby providing results that are in-line with (Hansen et al., 2024). We can measure any multicalibration metric with respect to either of two sets of groups. Prespecified Groups: refers to a small (less than $15$ across all datasets) set of protected groups that are formed using various features of the datasets. For this, we use the groups specifications that were used in  Hansen et al. (2024). Unspecified Groups: covers the case where the user does not specify which groups to protect against. In that case, the scores are expected to be multicalibrated with respect to all possible groups, and to calculate this, an extensive set of combinatorially generated groups are used.
@@
MCGrad uses LightGBM as GBDT implementation. To ensure seamless adoption by ML engineers, we set default hyperparameters so that MCGrad works out-of-the-box, without requiring application-specific tuning in the majority of use-cases. We determined the default hyperparameters by performing a grid search over $35$ company-internal datasets and selecting the configuration that never degraded the base model’s log loss or PRAUC, and minimized the average MCE across datasets. The resulting hyperparameters are: learning_rate=0.02873=0.02873, max_depth=5=5, min_child_samples =160=160, n_estimators=94=94, num_leaves=5=5, lambda_l2=0.00913=0.00913, min_gain_to_split=0.15=0.15. All other parameters are set as for LightGBM defaults.
For HKRR we follow Hansen et al. (2024) and pick the best hyperparameters from a set of four specifications using a held-out validation set. For DFMC we use the default hyperparameters of LightGBM since the paper does not suggest any specific hyperparameters.